# Towards Passive Safe Reinforcement Learning: A Comparative Study on Contact-rich Robotic Manipulation

### v1

@@ Abstract
Reinforcement learning (RL) has achieved remarkable success in various robotic tasks; however, its deployment in real-world scenarios, particularly in contact-rich environments, often overlooks critical safety and stability aspects. Policies without passivity guarantees can result in system instability, posing risks to robots, their environments, and human operators. In this work, we investigate the limitations of traditional RL policies when deployed in contact-rich tasks and explore the combination of energy-based passive control with safe RL in both training and deployment to answer these challenges. Firstly, we introduce energy-based constraints in our safe RL formulation to train passivity-aware RL agents. Secondly, we add a passivity filter on the agent output for passivity-ensured control during deployment. We conduct comparative studies on a contact-rich robotic maze exploration task, evaluating the effects of learning passivity-aware policies and the importance of passivity-ensured control. The experiments demonstrate that a passivity-agnostic RL policy easily violates energy constraints in deployment, even though it achieves high task completion in training. The results show that our proposed approach guarantees control stability through passivity filtering and improves the energy efficiency through passivity-aware training. A video of real-world experiments is available as supplementary material. We also release the checkpoint model and offline data for pre-training at Hugging Face
@@ I INTRODUCTION
In recent years, RL has earned increasing attention and success in addressing complex decision-making and control problems, especially in robotic applications [1]. From manipulation tasks to autonomous navigation, RL offers the potential to achieve unprecedented performance by learning optimal control policies. However, while RL excels in simulated environments, deploying these policies in real-world robotic systems remains a significant challenge [2]. This is particularly important in contact-rich tasks. Contact-rich tasks require the robots to interact with their environment in a safe and stable manner. Safety, particularly in terms of contact force regulation, is necessary to avoid damaging the environment, the robot, or the objects involved in the task. Stability, on the other hand, is crucial for maintaining predictable and controllable robot behavior, especially when interacting with complex and unpredictable surroundings [3]. For instance, an RL policy trained without considering system stability might generate control commands that cause high-frequency vibrations, or unbounded energy accumulation. Such behaviors can damage the robot, compromise the task outcome or pose safety risks to humans. Despite these risks, a substantial portion of RL research focuses primarily on maximizing task success rates or improving learning efficiency, with limited attention given to safety and stability. This gap presents a significant barrier to real-world deployment, where stability and safety are not only desirable but essential [4].
Traditional RL policies are often trained and tested in simulated environments without considering stability. This can not be ignored when deployed in real-world robotic tasks [1] since uncertainties and risks in realistic applications often cannot be eliminated. Moreover, contact-rich tasks amplify this risk due to the dynamic and sensitive nature of the interactions involved. Hence, ensuring that RL policies are inherently stable and safe during training and deployment is not only a theoretical concern but a practical necessity for robotics.
The safe RL concept has emerged to answer the safety concerns about the RL policies. Safe RL methods specify the safety requirements usually as constraints and aim to satisfy them while also solving the task [2]. The system can learn to avoid constraints by learning to predict the risk of states and actions from data [5]. In contact-rich tasks, self-proven control concepts such as variable impedance control (VIC) are used to add compliance to the system. A recent work [6] proposed a VIC-based safe RL method that effectively solves the contact-rich maze exploration problem. However, that work focused only on contact-related safety criteria, ignoring the control stability. In this work, we address this limitation by introducing passivity-based control.
Passivity is a fundamental concept in control theory. As proven in [7], passivity-based control is needed to facilitate physical interactions with the environment. It is vital to monitor and restrict the amount of energy allowed to be stored and injected by the controller into the robot to ensure the passivity and safety of the system [8]. Energy tanks are widely used to achieve passivity-based control by restricting the amount of energy allowed to be stored in the system [9, 10].
In the RL literature, the energy tanks were mostly limited to a deployment-time filtering step [11]. To the best of our knowledge, only [12] employed the classical energy-tank based passivity approach during RL training for contact-rich tasks by filtering the RL output and terminating the training episode when the tank is depleted. We advance the prior work by exploring passivity-aware learning by explicitly defining passivity-related constraints as part of the RL problem formulation. Leveraging a safe RL framework, we simultaneously answer the contact-rich safety and passivity problems. Furthermore, as stated in [13], solely limiting the total amount of energy is insufficient to ensure the passivity of the system. Hence, we augment the energy tank with a flow limit, unlike [12].
In summary, it can be concluded that while current safe RL methods can be utilized for learning control parameters of a VIC by maximizing rewards, the lack of the inclusion of passivity-based control formulations during deployment makes them ill-suited for contact-rich tasks. However, designing an augmented energy tank in terms of size and flow limits for contact-rich tasks is not trivial since the severity and frequency of unplanned interaction can affect energy spending, resulting in a premature depletion of the energy tank at which the task is entirely lost. Including such a tank in the training of the RL agent could lead to an improvement in the design process of augmented energy tanks as well as the overall performance during the deployment phase. Therefore, the presented work focuses on an augmented energy tank-based passivity formalism during the RL training phase to tackle the aforementioned challenges. More specifically, the formalism takes basis from the Safe-RL approach presented in  [6] and an augmented global energy tank-based approach inspired by [13] to ensure the passivity of the system. Furthermore, the presented work investigates the effect of such passivity-based formalism on the training/learning phase of the policy. Moreover, the performance in terms of task completion and energy expenditure of the different policies is compared to each other during deployment in simulation and the real world.
@@
    Investigation of RL policy stability: We analyze the limitation of traditional RL policies that lack passivity guarantees, especially deployed in complex scenarios, demonstrating their poor performance in contact-rich tasks.
@@
    Extensive study of different augmented energy-tank-based designs’ effects on the training of a RL policy: We introduce energy-tank-based constraints together with a contact force constraint to train passive safe RL policy.
@@
    Through extensive experiments both in simulation and the real world, we validate the effectiveness of our passivity-aware RL policy and passivity-ensured controller in contact-rich robotic tasks.
II Related work
II-A Passivity in robot compliant control
In order to facilitate physical interactions with the environment, it was formally proven in [7] that passivity-based control is needed. This becomes even more critical for kinematically redundant robots, as internal motions within the null space of the task Jacobian can result in an unintended increase in system energy, potentially leading to instability. Moreover, it was highlighted in [8] that it is vital to monitor and restrict the amount of energy allowed to be stored and injected by the controller into the robot to ensure the passivity and safety of the system. One used approach to guarantee passivity are energy tanks. However, as mentioned above and pointed out in [13], the bounding of the energy alone is not sufficient to ensure the passivity of the system, but also the energy flow between the controller and robot has to be limited. In recent years, works such as [14, 9, 10] have focused on restricting the energy flow with the energy tanks.
II-B Safe RL for contact-rich robotic tasks
RL can solve challenging contact-rich tasks [15]. In contact-rich tasks, VIC has proven to be an effective solution [16]. Reinforcement Learning (RL) can be leveraged to learn adaptive and safe VIC policies, enabling robust performance in more complex contact-rich scenarios. Specifically, some approaches [17, 18, 19] incorporate VIC parameters directly into the action space by utilizing demonstrations. Using Cartesian VIC as the RL action-space is shown to be effective for contact-rich tasks [20]. However, these works do not explicitly address the safety issues. The safety can be formulated by introducing constraints to the RL policy in the form of a constrained Markov decision process (CMDP) [21]. A way to solve CMDPs is by training a safety critic to predict a potential risk value before taking action and train a conservative policy [22]. [5] proposed a recovery mechanism by training a recovery policy to enhance safety. Instead of collision-avoidance as in [5], a recent work [6] integrates a safety critic for contact-rich tasks, by focusing on how to stay safe during and after a collision without sacrificing task completion. However, safe RL policies for contact-rich tasks rarely consider the system’s stability, which is essential in real-world deployment. We take a step towards a passive safe RL policy for robotic contact-rich tasks that achieve dual safe and stable performance.
@@ II-B Safe RL for contact-rich robotic tasks
Figure 1: System overview. The passivity-agnostic safe RL framework (a) constitutes the baseline method. We propose a passivity-aware policy (b) through energy constraints $C_{e}$, and a passivity-ensured system (c) by adding a passivity layer to stabilize the RL action. We evaluate these three approaches through simulation and real-world experiments (d). Commonly, in all approaches, the RL agent observes the environment and outputs an action applied to the Cartesian VIC controller (Sec. III-B). The safe RL agent (e) uses a safety critic model to detect high risk and switch to the safer recovery policy. These models are pre-trained using offline data (f) consisting of random actions executed in multiple areas of the maze. (Sec. IV-B1). The policy training (g) is completed by online training in simulation (Sec. IV-B2).
II-C Passivity considerations in RL
A passive RL policy can provide benefits in terms of stability and task performance. Work such as [11, 23] uses an RL-based method to train a passive policy; however, the safety issue was not addressed, and the passivity guarantee is more complex in model-agnostic environment [23]. Zanella et al. [12] use an energy tank-based filter to enforce passivity in learned policies. However, their approach is validated only in simulation while ensuring passive performance in real-world deployment remains crucial. In addition, the utilization of an energy tank with the augmented flow limit at its port was not investigated in  [12], which is also necessary to bind the energy flow between the controller and robot as mentioned in [13]. Different from the previous research, we go further based on a Safe RL method by not only extending it with passivity-aware RL policy training but also with passivity-ensured deployment in real-world contact-rich tasks.
III Passive Safe Reinforcement Learning
@@
The presented approach draws upon passive control and safe reinforcement learning fields to learn safe and stable policies for solving contact-rich tasks. As previously mentioned, it uses a passivity-based approach to observe and limit the energy stored in the system such that it can both: 1. impose constraints on the safe RL to train passivity-aware policies (Sec. III-D), and 2. achieve passivity-ensured control by a passivity layer running on the policy output (Sec. III-E). The former approach by itself does not guarantee passivity due to the incompleteness of data and the stochasticity of learning methods. The latter approach provides this guarantee, but can significantly reduce task performance by disrupting the task-oriented policy behavior. We combine both approaches to achieve good task performance while ensuring stability. A diagram of our work is shown in Fig. 1.
III-A Safe RL for contact-rich tasks
@@
This section briefly introduces the contact-rich safe RL framework [6] that forms the basis of our proposed method. The problem is formulated as a CMDP [21]. A CMDP can be expressed as a tuple $(\mathcal{S},\mathcal{A},R,P,\gamma,\mu,\mathcal{C})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $P$ is the state transition probability, $\gamma\in(0,1)$ is the reward discount factor, $\mu$ is the starting state distribution and $\mathcal{C}=\left\{\left(c_{i}:\mathcal{S}\rightarrow\{0,1\},\chi_{i}\in\mathbb{R}\right)\mid i\in\mathbb{Z}\right\}$ denotes the constraints that the agent must satisfy. For this specific formalism, we define the state space, action space, reward function, and safety constraints as below. All vectors are in the world reference frame unless specified.
@@
  • State Space $\mathcal{S}$. The state space is designed as A 6-dimensional vector for safety critic network and recovery policy including the wrench $[f_{x},f_{y},f_{z},m_{x},m_{y},m_{z}]\in se^{*}(3)$ measured from an F/T sensor. Additionally, a 9-dimensional state vector is used, which extends this representation by incorporating the end-effector position [$p_{x},p_{y},p_{z}$$\in\mathbb{R}$.
@@
  • Action Space $\mathcal{A}$. Based on  [6], we design a 4-dimensional vector as action space, consisting of two end-effector displacement values $[\Delta p_{x},\Delta p_{y}]\in\mathbb{R}$ and 2 stiffness parameters $[K_{1},K_{2}]$. The action dimensions are bounded as listed in Table I. We keep the position $\Delta p_{z}$ and stiffness $K_{z}$ in the z-axis constant. The stiffness values are defined w.r.t. the motion vector $\Delta\mathbf{p}$ with $K_{1}$ along the motion vector, $K_{2}$ orthogonal to the motion vector and parallel to the ground plane, and $K_{z}$ normal to the ground plane.
@@
  • Reward Function $R$. The agent is penalized for the Euclidean distance between the current position of the end-effector and the goal position multiplied by a constant $r_{pos}$. High-force collisions are penalized with $r_{col}$. A bonus is given for reaching the goal $r_{goal}$ to increase convergence speed. The reward multipliers used in our experiments are presented in Table I.
@@
  • Constraints $\mathcal{C}$. The force constraint $c_{f}$ requires the measured force magnitude to stay below a threshold. We also propose two new constraints for passive control, described later in Sec. III-D.
The Safe RL network structure is shown in Fig. 1.e, there are three components in a safe RL agent: Firstly, a safety critic model and recovery policy are trained with the pre-collected offline data. Then, task policy is trained online with the maze exploration task. The policy output as action is given to the VIC, which controls the robot to achieve the desired actions. In the following section, we describe how we apply the RL action with our low-level controller.
III-B Cartesian impedance control
@@
$\boldsymbol{\tau}^{\top}_{\text{IC}}=\boldsymbol{\tau}^{\top}_{\text{K}}-\boldsymbol{\tau}^{\top}_{\text{D}},$ (1)
@@
${\mathbf{w}^{\text{0},\text{EE}}_{K}}^{\top}=\begin{bmatrix}\mathbf{K}_{t}&\mathbf{K}_{c}\\ \mathbf{K}^{\top}_{c}&\mathbf{K}_{r}\\ \end{bmatrix}\Delta{\boldsymbol{\eta}}.$ (2)
Here, the infinitesimal body twist displacement is described by $\Delta{\boldsymbol{\eta}}\in se(3)$. The positive definite matrice $\mathbf{K}_{c}\in\mathbb{R}^{3\times 3}$ describes the decoupling between the stiffness values of the rotational ($\mathbf{K}_{r}\in\mathbb{R}^{3\times 3}$), and translational ($\mathbf{K}_{t}\in\mathbb{R}^{3\times 3}$) axes. In this work the values of translational stiffness matrix $\mathbf{K}_{t}$ are obtained by mapping the RL action stiffness matrix $\mathbf{K}_{a}=\operatorname{diag}(K_{1},K_{2},K_{z})$ (Sec. III-A) to the world frame as follows $\mathbf{K}_{t}=\mathbf{R}_{p}^{\top}\mathbf{K}_{a}\mathbf{R}_{p}$, given
@@
III-C Passivity-based approach for variable stiffness control

### v2

# Passivity-Centric Safe Reinforcement Learning for Contact-Rich Robotic Tasks

@@ Abstract
Reinforcement learning (RL) has achieved remarkable success in various robotic tasks; however, its deployment in real-world scenarios, particularly in contact-rich environments, often overlooks critical safety and stability aspects. Policies without passivity guarantees can result in system instability, posing risks to robots, their environments, and human operators. In this work, we investigate the limitations of traditional RL policies when deployed in contact-rich tasks and explore the combination of energy-based passive control with safe RL in both training and deployment to answer these challenges. Firstly, we reveal the discovery that standard RL policy does not satisfy stability in contact-rich scenarios. Secondly, we introduce a passivity-aware RL policy training with energy-based constraints in our safe RL formulation. Lastly, a passivity filter is exerted on the policy output for passivity-ensured control during deployment. We conduct comparative studies on a contact-rich robotic maze exploration task, evaluating the effects of learning passivity-aware policies and the importance of passivity-ensured control. The experiments demonstrate that a passivity-agnostic RL policy easily violates energy constraints in deployment, even though it achieves high task completion in training. The results show that our proposed approach guarantees control stability through passivity filtering and improves the energy efficiency through passivity-aware training. A video of real-world experiments is available as supplementary material. We also release the checkpoint model and offline data for pre-training at Hugging Face.
@@ I INTRODUCTION
In recent years, RL has earned increasing attention and success in addressing complex decision-making and control problems, especially in robotic applications [1]. From manipulation tasks to autonomous navigation, RL offers the potential to achieve unprecedented performance by learning optimal control policies. However, while RL excels in simulated environments, deploying these policies in real-world robotic systems remains a significant challenge due to the safety concerns [2]. This is particularly important in contact-rich tasks. Contact-rich tasks require the robots to interact with their environment safely, particularly in terms of contact force regulation, which is necessary to avoid damaging the environment, the robot, or the objects involved in the task. Although, there has been a significant effort to answer the contact-rich safety problem, the role of the control stability is often neglected within RL.
Stability is crucial for maintaining predictable and controllable robot behavior, especially when interacting with complex and unstructured surroundings [3] which can not be ignored in real-world robotic deployment [1]. For instance, an RL policy trained without considering system stability might generate control commands that cause high-frequency vibrations or unbounded energy accumulation. Such behaviors can damage the robot, compromise its performance, or pose safety risks to humans. Hence, ensuring that RL policies are inherently stable and safe is not only a theoretical concern but a practical necessity for robotics. Despite these risks, a substantial portion of RL research focuses primarily on maximizing task success rates or improving learning efficiency, with limited attention given to the stability of policy. This gap presents a significant barrier to real-world deployment, where stability and safety are not only desirable but essential [4].
Passivity-based control (PBC) is a fundamental method to prove the stability of physical interactions with the environment [5]. It is even more critical for kinematically redundant robots, as internal motions within the null space of the task Jacobian can result in an unintended increase in system energy, potentially leading to instability. Moreover, it is vital to monitor and restrict the amount of energy allowed to be stored and injected by the controller into the robot to ensure the passivity and safety of the system [6]. Energy tanks are widely used to achieve passivity-based control by restricting the amount of energy allowed to be stored in the system [7]. However, as mentioned above and pointed out in [8], the bounding of the energy alone is not sufficient to ensure the passivity of the system, but also the energy flow between the controller and robot has to be limited. We follow the recent works [9, 7, 10] with an augmented energy tank approach that also restricts the energy flow. Although energy tank-based passivity is studied extensively in the classical control literature, its inclusion in RL-based control learning is not sufficiently explored.
The inclusion of PBC in an RL system constitutes a significant gap in the RL literature. In this work, we aim to explore different modes of combination between RL and PBC, and study their implications. We term the standard RL approaches that do not include any passivity treatment as passivity-agnostic RL (Fig. 1.a). Then, we distinguish between the passivity-filtered approach, that imposes a passivity filter at the RL output; and the passivity-aware approach which aims to learn the passivity in the RL policy. In the former, the RL policy is trained traditionally, however, the filter analytically calculates the system energy and modifies the RL actions to stay passive in deployment. While it guarantees the passivity of the control, the RL agent remains unaware of the passivity. The latter introduces the energy awareness in the policy, however, it cannot guarantee the passivity as it is a black-box policy. Subsequently, we propose the passivity-ensured RL (Fig. 1.c) that includes both passivity guarantee and energy efficiency by combining the two approaches, as summarized in Table I.
Passivity-aware learning offers the possibility to obtain policies that solve tasks while providing benefits in terms of energy economy and long-term performance. Additionally, for systems with complex, nonlinear dynamics where analytical PBC design becomes intractable, learning provides a practical alternative that can discover passive policies without requiring complete system modeling. Work such as [11] and [12] use RL to enhance designed PBC systems. [11] uses RL to learn the parameters of the control system. [12] uses RL to dynamically adapt the size of an energy tank for changing tasks. To the best of our knowledge, only [13] employs the classical energy-tank based passivity approach during RL training by filtering the RL output and terminating the training episode when the energy tank is depleted. Although they train the system with an active filter, the RL agent is expected to learn the passivity only indirectly.
TABLE I: Method comparison
passivity- agnostic passivity- filtered passivity- aware passivity- ensured passivity guarantee energy efficiency
We advance the prior work by explicitly defining passivity-related constraints as part of the RL problem formulation. Furthermore, as stated in [8], solely limiting the total amount of energy is insufficient to ensure the passivity of the system. Hence, we augment the energy tank with a flow limit, unlike [13]. The main impact of this work lies in a systematic comparison of different approaches to combine PBC and RL. Furthermore, we formulate our passive-aware learning approach on a state-of-the-art safe RL framework to explicitly answer the contact-rich safety and passivity problems together. The safe RL concept has emerged to answer the safety concerns about the RL policies, which specify the safety requirements usually as constraints and aim to satisfy them while also solving the task [2]. We adopt a recent safe RL method that combines variable impedance control (VIC) and safety critic frameworks for contact rich tasks [14]. Leveraging this safe RL framework, we take a step towards a passivity-ensured safe RL policy for robotic contact-rich tasks that achieves dual safe and stable performance.
Lastly, we compare the four approaches of passivity-agnostic, passivity-filtered, passivity-aware and passivity-ensured RL extensively on a contact-rich interaction problem. The maze exploration task requires the robot to find the exit of a path by blindly touching the walls. Being contact-rich indicates a long exposure to physical contact. In fact, in this task, the contact is not only inevitable, but also fundamental to solve the task as the vision modality is not available. Our experimental study investigates the effect of passivity-based formalisms on the training/learning phase of the policy. Moreover, the performance in terms of task completion and energy expenditure of the different policies is compared to each other during deployment in simulation and the real world. The results confirm the advantages of the proposed approach in both guaranteeing the passivity and improving the energy usage while also safely solving a contact-rich task.
@@
    Extensive study of different RL-based passive control learning approaches (Table I), and their effect on both training and deployment phases, evaluated from safety, stability, energy efficiency and task completion perspectives.
@@
    Design and evaluation of the augmented energy tank-based passivity formulation in RL, with varying energy tank and energy flow constraints.
@@
    Introduction of passivity-ensured safe RL framework that simultaneously answers both control passivity and contact-rich safety problems.
@@
Figure 1: System overview. The passivity-agnostic safe RL framework (a) constitutes the baseline method. Our contribution includes passivity-aware training (b) through energy constraints $C_{e}$, and passivity-ensured deployment (c) by adding a passivity layer to stabilize the RL action. We evaluate these three approaches through simulation and real-world experiments (d). Commonly, in all approaches, the RL agent observes the environment and outputs an action applied to the Cartesian VIC controller (Sec. II-B). The safe RL agent (e) uses a safety critic model to detect high risk and switch to the safer recovery policy. These models are pre-trained using offline data (f) consisting of random actions executed in multiple areas of the maze. The policy training (g) is completed by online training in simulation.
II Passive Safe Reinforcement Learning
@@
The presented approach draws upon passive control and safe reinforcement learning fields to learn safe and stable policies for solving contact-rich tasks. As previously mentioned, it uses a passivity-based approach to observe and limit the energy stored in the system such that it can both: 1. impose constraints on the safe RL to train passivity-aware policies (Sec. II-D), and 2. achieve passivity-ensured control by a passivity layer running on the policy output (Sec. II-E). The former approach by itself does not guarantee passivity due to the incompleteness of data and the stochasticity of learning methods. The latter approach provides this guarantee, but can significantly reduce task performance by disrupting the task-oriented policy behavior. We combine both approaches to achieve good task performance while ensuring stability. A diagram of our work is shown in Fig. 1.
II-A Safe RL for contact-rich tasks
This section briefly introduces the contact-rich safe RL framework that forms the basis of our proposed method. The Safe RL network structure is shown in Fig. 1.e. Our RL agent consists of different components such as the safety critic model and recovery policy, in addition to the classic RL task policy. The safety critic is a separate network that predicts the potential risk of state-action pairs [15]. In the case of high risk, a recovery policy is activated to escape from risky areas [16]. The safety critic model and recovery policy are trained with the pre-collected offline data focusing only on the constraint violations. Then, all components including task policy are trained online with the maze exploration task. The policy output as action is given to the VIC, which controls the robot to achieve the desired actions. We use a variable impedance action-space for its advantages in physical interaction tasks [14, 17].
@@ II-A Safe RL for contact-rich tasks
Safety can be formulated by introducing constraints to the RL policy in the form of a constrained Markov decision process (CMDP) [18]. A CMDP is expressed as a tuple $(\mathcal{S},\mathcal{A},R,P,\gamma,\mu,\mathcal{C})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $P$ is the state transition probability, $\gamma\in(0,1)$ is the reward discount factor, $\mu$ is the starting state distribution and $\mathcal{C}=\left\{\left(c_{i}:\mathcal{S}\rightarrow\{0,1\},\chi_{i}\in\mathbb{R}\right)\mid i\in\mathbb{Z}\right\}$ denotes the constraints that the agent must satisfy. For this specific formalism, we define the state space, action space, reward function, and safety constraints as below. All vectors are in the world reference frame unless specified.
@@
  • State Space $\mathcal{S}$. The state space is designed as A 6-dimensional vector for safety critic network and recovery policy including the wrench $[f_{x},f_{y},f_{z},m_{x},m_{y},m_{z}]\in se^{*}(3)$ measured from an F/T sensor. Additionally, a 9-dimensional state vector is used, which extends this representation by incorporating the end-effector position [$p_{x},p_{y},p_{z}$]$\in\mathbb{R}^{3}$.
@@
  • Action Space $\mathcal{A}$. Based on  [14], we design a 4-dimensional vector as action space, consisting of two end-effector displacement values $[\Delta p_{x},\Delta p_{y}]\in\mathbb{R}^{2}$ and 2 stiffness parameters $[k_{1},k_{2}]$. The action dimensions are bounded as listed in Table II. We keep the position $\Delta p_{z}$ and stiffness $k_{z}\in\mathbb{R}$ in the z-axis constant. The stiffness values are defined w.r.t. the motion vector $\Delta\mathbf{p}$ with $k_{1}\in\mathbb{R}$ along the motion vector, $k_{2}\in\mathbb{R}$ orthogonal to the motion vector and parallel to the ground plane, and $k_{z}$ normal to the ground plane.
@@
  • Reward Function $R$. The agent is penalized for the Euclidean distance between the current position of the end-effector and the goal position multiplied by a constant $r_{pos}$. High-force collisions are penalized with $r_{col}$. A bonus is given for reaching the goal $r_{goal}$ to increase convergence speed. The reward multipliers used in our experiments are presented in Table II.
@@
  • Constraints $\mathcal{C}$. The force constraint $c_{f}$ requires the measured force magnitude to stay below a threshold. We also propose two new constraints for passive control, described later in Sec. II-D.
In the following section, we describe how we apply the RL action with our low-level controller.
II-B Cartesian impedance control
@@
$\boldsymbol{\tau}^{\top}_{\text{IC}}=\boldsymbol{\tau}^{\top}_{\text{K}}-\boldsymbol{\tau}^{\top}_{\text{D}},\vspace{-0.1cm}$ (1)
@@
${\mathbf{w}^{\text{0},\text{EE}}_{K}}^{\top}=\begin{bmatrix}\mathbf{K}_{t}&\mathbf{K}_{c}\\ \mathbf{K}^{\top}_{c}&\mathbf{K}_{r}\\ \end{bmatrix}\Delta{\boldsymbol{\eta}}.\vspace{-0.05cm}$ (2)
Here, the infinitesimal body twist displacement is described by $\Delta{\boldsymbol{\eta}}\in se(3)$. The positive definite matrix $\mathbf{K}_{c}\in\mathbb{R}^{3\times 3}$ describes the decoupling between the stiffness values of the rotational ($\mathbf{K}_{r}\in\mathbb{R}^{3\times 3}$), and translational ($\mathbf{K}_{t}\in\mathbb{R}^{3\times 3}$) axes. In this work the values of translational stiffness matrix $\mathbf{K}_{t}$ are obtained by mapping the RL action stiffness matrix $\mathbf{K}_{a}=\operatorname{diag}(k_{1},k_{2},k_{z})$ (Sec. II-A) to the world frame as follows $\mathbf{K}_{t}=\mathbf{R}_{p}^{\top}\mathbf{K}_{a}\mathbf{R}_{p}$, given
@@
II-C Passivity-based approach for variable stiffness control